Information from the financial statements of Henderson-Niles Industries included the following at December 31, 2021:

Common shares outstanding throughout the year 100 million
Convertible preferred shares (convertible into 36 million shares of common) 60 million
Convertible 8% bonds (convertible into 14.5 million shares of common) $ 1,600 million


Henderson-Niles’s net income for the year ended December 31, 2021, is $740 million. The income tax rate is 25%. Henderson-Niles paid dividends of $3 per share on its preferred stock during 2021.

Required:
Compute basic and diluted earnings per share for the year ended December 31, 2021. (Enter your answers in millions (i.e., 10,000,000 should be entered as 10). Round "Earnings per share" answers to 2 decimal places.)

Numerator / Denominator = Earnings per share
Basic $560selected answer correct / not attempted = not attempted
Diluted $560selected answer incorrect / not attempted = not attempted
Answer #1

Basic earnings per sharing = (Net income - Preferred dividend)/No. Of outstanding common stock

= (740-60*3)/100 = $5.6 per share

Diluted earnings per share = (Net income + Saving of interest net of tax)/Total potential common stock

= (740+1600*8%*75%)/(100+36+14.5)

= $5.55 per share